OpenStudy (anonymous):

Its just a simple inequality question: 12 / (x-3) > (x+1) But I have some problem not getting the right answer.

OpenStudy (anonymous):

\[12>(x-3)(x+1)\] Then, \[12>x ^{2}-2x-3\] \[0>x ^{2}-2x-15\]

OpenStudy (saifoo.khan):

@Rizki_Zhou , are you sure?

OpenStudy (anonymous):

we can cange position, \[x ^{2}-2x-15<0\]

OpenStudy (anonymous):

@saifoo.khan yes, is it false?

OpenStudy (saifoo.khan):

@Rizki_Zhou , yea.. idk im getting a feeling.

OpenStudy (anonymous):

@saifoo.khan But I think i do it well.

OpenStudy (saifoo.khan):

\[\frac{12}{x-3} - x-1 > 0\] \[\frac{12 +(-x-1)(x-3)}{x-3} > 0\] \[\frac{-x^2 + 2x + 3 + 12}{x-3} > 0\] \[\frac{-x^2 + 2x + 15}{x-3} > 0\]

OpenStudy (saifoo.khan):

x^2 - 2x - 15 < 0 (x-5)(x+3) <0 x < 5 and x < -3 Therefore, -3<x<5

OpenStudy (saifoo.khan):

Dont forget about the denominator, \[\frac{1}{x-3} > 0\]\[x - 3 < 0\]\[x < 3\]

OpenStudy (saifoo.khan):

Now, the two inequalities are, -3<x<5, x < 3

OpenStudy (anonymous):

okay @saifoo.khan, got it! thank you :)

OpenStudy (saifoo.khan):

Welcome. =D